OpenStudy (anonymous):

Find the center of this ellipse: 9x^(2)+y^(2)+36x-4y+31=0

OpenStudy (anonymous):

@lgbasallote or @satellite73 you know these?

OpenStudy (lgbasallote):

not good with graphs sorry

OpenStudy (anonymous):

these are the worst references ever

OpenStudy (anonymous):

9x^(2)+y^(2)+36x-4y+31=0 or, 9 (x^2+4x +4) +(y^2-4y+4)-9=0 or, 9(x+2)^2 + (y-2)^2 = 9 or, (x+2)^2 +(y-2)^2 /9 = 1.... see that the centre is (-2,2)

OpenStudy (anonymous):

The purplemath.com reference is exactly what you need to do this.

OpenStudy (zzr0ck3r):

what is wrong with @bhaweshwebmasters answer?

OpenStudy (zzr0ck3r):

do you know how to complete the square? Thats all we are doing on these.